OpenStudy (anonymous):

IF THERE ARE 247,947 PEOPLE STUDYING THIS SUBJECT, WHY WONT ANYONE HELP!!

OpenStudy (anonymous):

What's the question?

OpenStudy (alexandervonhumboldt2):

worng: only 308 people online now

OpenStudy (thread252):

the website lied, we are all computers!!!!!!! LOL

OpenStudy (anonymous):

A

OpenStudy (theedwardsfamily):

Because everyone's busy eating cookies and #netflixandchill xDDD

OpenStudy (carlyleukhardt):

I AGREE WITH YOU ITS SO ANNOYING!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

When adding 6.4 + 2.85 (10^11)

Vocaloid (vocaloid):

you just have to add 6.4 + 2.85 and keep the 10^11 part at the end

OpenStudy (anonymous):

OH

OpenStudy (theedwardsfamily):

use symbolab c:

OpenStudy (anonymous):

what is symbolab?

OpenStudy (theedwardsfamily):

https://www.symbolab.com/

OpenStudy (anonymous):

I DONT WANT A COUCULATOR I WANT AN EXPLINATION GEEZ I WANT TO USE MY BRAIN NOT WAST IT!!

OpenStudy (anonymous):

When adding 6.4 + 2.85 (10^11) you just have to add 6.4 + 2.85 and keep the 10^11 part at the end (These are the explanations - they both were mentioned above)

OpenStudy (theedwardsfamily):

it explains @MariahDimond -_-

OpenStudy (anonymous):

DONT CARE I DONT UNDERSTAND IT @TheEdwardsFamily

OpenStudy (anonymous):

@Zale101 @livias.random its 9.25

OpenStudy (theedwardsfamily):

ok ok fine then. ._.

OpenStudy (zale101):

Explain the process of how you got that answer.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i added 6.4 + 2.85 \[6.4+2.85=9.25 exponents --> adding 11+11=22\]

OpenStudy (anonymous):

so its \[9.25^{22}\]

OpenStudy (zale101):

Correct. Yo were right about adding the coefficients 6.4 and 2.85. However, you had \(\large 6.4*10^{11}+2.85*10^{11}\) Both of these terms have \(10^{11}\). We do no add exponents but instead we keep it the same.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so that means its \[6.4*10^{11} + 2.85 * 10^{11}= 9.25*10^{11} :)\]

OpenStudy (zale101):

Yes.

OpenStudy (anonymous):

is it ok if you help me with more ? Mr.zale?

OpenStudy (anonymous):

That next one has a division. When you divide 10^a by 10^b to find the new power you do 10^(a-b)

OpenStudy (anonymous):

ok but i still dont get what do i divide?

OpenStudy (anonymous):

Ms.livias?

OpenStudy (anonymous):

Hi! Sorry about that... So the integer in front of the 10 (in this case, 2 and 8) is divided as normal. So 2÷8=0.25

OpenStudy (anonymous):

so then i divide the eponents

OpenStudy (anonymous):

or multiply?

OpenStudy (anonymous):

then you subtract the exponent on the bottom from the one on top (In this case, 21-3)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its 18

OpenStudy (anonymous):

\[0.25*10^{18}\]

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

is it ok if you help me with more?

OpenStudy (anonymous):

I can't be around for much longer sorry...

OpenStudy (samsungfanboy):

@TheEdwardsFamily lel ;)

OpenStudy (theedwardsfamily):

i know... lol

OpenStudy (anonymous):

what?

OpenStudy (theedwardsfamily):

@SamsungFanBoy ;)